Ultraviolet light of wavelength Lambda one and Lambda to when allowed to fall on hydrogen atom in the ground state is found to liberate electrons with kinetic energy 1.8 electron volt and 4 electron volt respectively find the value of Lambda one / Lambda two (L1/L2)

Grade:12

1 Answers

Aakash
6 years ago
By De Broglie equation; λ= h/mc which can also be written as λ = h/√2mK.E => λ1 = h/√2×m×1.8 λ2 = h/√2×m×4λ1/λ2 = √4/√1.8 = √2.2 = 1.48 Hence, the ratio of λ1 and λ2 is 1.48:1 or 3:2